# Topology-Informed Graph Transformer

### v2

@@ Appendix A Mathematical Proofs
This subsection focuses on listing the mathematical background required for proving a series of theorems outlined in the main text of the paper. Throughout this subsection, we regard a graph $G:=(V,E)$ as a 1-dimensional topological space endowed with the closure-finiteness weak topology (CW topology), the details of which are written in (Hatcher, 2002)[Chapter 0, Appendix]. In particular, we may regard $G$ as a 1-dimensional CW complex, the set of nodes of $G$ corresponds to the 00-skeleton of $G$, and the set of edges of $G$ corresponds to the $1$-skeleton of $G$.
@@
#$\#E(G)-\#V(G)=\#\text{ Connected components of }G-\#\text{ cycle basis of }G$
@@
[…]
=#$=\#E(H\setminus\phi(\{v\}))-\#V(H\setminus\phi(\{v\}))$
=#$=\#E(H)-\#V(H)+1$
=#$=\#E(G)-\#V(G)+1$
=#$=\#E(G\setminus\{v\})-\#V(G\setminus\{v\})$
[…]
@@
The geometric rate of convergence $\gamma$ satisfies the inequality $10<\gamma<$ 1. We note that the value of $\gamma$ is determined from eigenvalues of the matrix $N:=D^{-1/2}\textbf{M}D^{1/2}$, all of whose eigenvalues excluding the largest eigenvalue is known to have absolute values between 00 and $1$ for non-bipartite graphs $G$ (Lovasz, 1993)[Section 3]. To obtain the statement of the theorem, we apply above equation with probability distributions $\delta$ whose $i$-th component is $1$, and all other components are equal to 00.
@@ Appendix B Abalation study
TIGT 0.057$\pm$0.002
w/o graph information 0.059$\pm$0.005
w/o topological positional embedding 0.060$\pm$0.003
not share weight in topological positional embedding 0.061$\pm$0.003
Transformer $\rightarrow$ Performer in global attention 0.063$\pm$0.001
w/o global attention 0.063$\pm$0.003
Tanh $\rightarrow$ ReLU in topological positional embedding 0.063$\pm$0.004
Dual-path MPNNs $\rightarrow$ Single-path MPNNs 0.064$\pm$0.003
Sum $\rightarrow$ Mean readout in graph information 0.069$\pm$0.001
Cy2C-GNNs 0.102$\pm$0.002
Cy2C-GNNs(Large) 0.121$\pm$0.003
GraphGPS 0.070$\pm$0.004
@@ C.1 Datasets
A detail of statistical properties of benchmark datasets are summarized in Table 5. We perform the experiments on GraphGPS (Rampášek et al., 2023) framework.
@@ C.2 Additional notes on model design
TIGT includes batch normalization layer, proposed from  (Ioffe & Szegedy, 2015), and residual connection, as outlined in  (He et al., 2016).
@@ C.3 Hyperparameters
To ensure a fair comparison, we followed the hyperparameter settings of GraphGPS (Rampášek et al., 2023) as outlined in their benchmark datasets. It’s worth noting that, due to the intrinsic nature of the TIGT architecture, the number of model parameters varies. Details regarding these hyperparameters are provided in Table 6.
@@ C.4 Implementation detail of Experiment on CSL dataset
The CSL dataset (Murphy et al., 2019) was obtained using the ’GNNBenchmarkDataset’ option from the PyTorch Geometric library (Fey & Lenssen, 2019). We partitioned the dataset into training, validation, and test sets with proportions of 0.6, 0.2, and 0.2, respectively. Detailed descriptions of the hyperparameters are presented in Table 7. Hyperparameters for the CSL dataset that are not specified here are consistent with those used in the ZINC dataset experiment (Rampášek et al., 2023; Ma et al., 2023).

### v3

@@ Appendix A Mathematical Proofs
This subsection focuses on listing the mathematical background required for proving a series of theorems outlined in the main text of the paper. Throughout this subsection, we regard a graph $G:=(V,E)$ as a 1-dimensional topological space endowed with the closure-finiteness weak topology (CW topology), the details of which are written in (Hatcher, 2002)[Chapter 0, Appendix]. In particular, we may regard $G$ as a 1-dimensional CW complex, the set of nodes of $G$ corresponds to the 0-skeleton of $G$, and the set of edges of $G$ corresponds to the $1$-skeleton of $G$.
@@
where the term ”# $\text{ cycle basis of }G$” is the number of elements of a cycle basis of $G$. This number is well-defined regardless of the choice of a cycle basis, because its number is equal to the dimension of the first homology group of $G$ with rational coefficients, one of the topological invariants of $G$.
@@
# $\;\;\;\;\;\#\text{ Connected components of }H\setminus\phi(\{v\})-\#\text{ cycle basis of }H\setminus\phi(\{v\})$
$=\#E(H\setminus\phi(\{v\}))-\#V(H\setminus\phi(\{v\}))$
$=\#E(H)-\#V(H)+1$
$=\#E(G)-\#V(G)+1$
$=\#E(G\setminus\{v\})-\#V(G\setminus\{v\})$
=# $=\#\text{ Connected components of }G\setminus\{v\}-\#\text{ cycle basis of }G\setminus\{v\}$
@@
The geometric rate of convergence $\gamma$ satisfies the inequality $0<\gamma<1$. We note that the value of $\gamma$ is determined from eigenvalues of the matrix $N:=D^{-1/2}\textbf{M}D^{1/2}$, all of whose eigenvalues excluding the largest eigenvalue is known to have absolute values between 0 and $1$ for non-bipartite graphs $G$ (Lovasz, 1993)[Section 3]. To obtain the statement of the theorem, we apply above equation with probability distributions $\delta$ whose $i$-th component is $1$, and all other components are equal to 0.
@@ Appendix B Abalation study
[…]
@@ C.1 Datasets
A detail of statistical properties of benchmark datasets are summarized in Table 5. We perform the experiments on GraphGPS (rampášek2023recipe) framework.
@@ C.2 Additional notes on model design
TIGT includes batch normalization layer, proposed from  (Ioffe and Szegedy, 2015), and residual connection, as outlined in  (He et al., 2016).
@@ C.3 Hyperparameters
To ensure a fair comparison, we followed the hyperparameter settings of GraphGPS (rampášek2023recipe) as outlined in their benchmark datasets. It’s worth noting that, due to the intrinsic nature of the TIGT architecture, the number of model parameters varies. Details regarding these hyperparameters are provided in Table 6.
@@ C.4 Implementation detail of Experiment on CSL dataset
The CSL dataset (Murphy et al., 2019) was obtained using the ’GNNBenchmarkDataset’ option from the PyTorch Geometric library (Fey and Lenssen, 2019). We partitioned the dataset into training, validation, and test sets with proportions of 0.6, 0.2, and 0.2, respectively. Detailed descriptions of the hyperparameters are presented in Table 7. Hyperparameters for the CSL dataset that are not specified here are consistent with those used in the ZINC dataset experiment (rampášek2023recipe; Ma et al., 2023).